The perimeter of a triangle is 69 cm. Side a is 5 cm shorter than side b. Side c is twice as long as side a. Find the length of each side. I need to know how to do it in 2 equations. Thanx!
Answer by stanbon(75887)

The perimeter of a triangle is 69 cm. Side a is 5 cm shorter than side b. Side c is twice as long as side a. Find the length of each side. I need to know how to do it in 2 equations.
----------------------------
Equations:
a = b-5
c = 2a
-------
a + b + c = 69
a + (a+5) + 2a = 69
4a + 5 = 69
4a = 64
a = 16
b = a+5 = 21
c = 2a = 32
-----------
Cheers,
Stan H.